# Surprisal and Metaphor Novelty: Moderate Correlations and Divergent Scaling Effects

### v1

@@ A.2 GPT4o-METANOV Construction
To generate the sentences in GPT4o-METANOV, we prompt GPT4o once for each word in the dataset. As we planned for 10 different words, we prompted the model 10 consecutive times to construct this dataset. ‘‘I am curating a dataset to be used in a study about metaphoric knowledge in pretrained language models PLMs e.g. GPT-2. The dataset should consist of sentences that correspond to target nouns/verbs. For each target noun/verb, there should be 10 sentences. The target noun/verb should be used 10 times using conventional metaphoric meanings of the target noun/verb, and 10 times using novel metaphoric meanings of the target noun/verb. Please suggest a target noun, and generate 20 sentences following the requirements above: 10 sentences with Conventional Metaphor usages, and 10 sentences with Novel Metaphor usages.’’
@@ B.1 Distributions
In Figure 3, we plot the distributions of novelty scores and surprisal of the VUA-METANOV.
@@
Figure 4: Scatter plots between metaphor novelty scores on the x-axis and surprisal on the y-axis at different relative positions of metaphoric words in sentences. 0-10% means metaphoric words located at the first 1/10Length1/10*Length of the sentence.

### v2

@@ A.2 GPT-4o-metaphors Construction
To generate the sentences in GPT-4o-metaphors, we prompt GPT-4o once for each word in the dataset. As we planned for 10 different words, we prompted the model 10 consecutive times to construct this dataset. ‘‘I am curating a dataset to be used in a study about metaphoric knowledge in pretrained language models PLMs e.g. GPT-2. The dataset should consist of sentences that correspond to target nouns/verbs. For each target noun/verb, there should be 10 sentences. The target noun/verb should be used 10 times using conventional metaphoric meanings of the target noun/verb, and 10 times using novel metaphoric meanings of the target noun/verb. Please suggest a target noun, and generate 20 sentences following the requirements above: 10 sentences with Conventional Metaphor usages, and 10 sentences with Novel Metaphor usages.’’
@@ B.1 Distributions
In Figure 3, we plot the distributions of novelty scores/labels and surprisal of the VUA datasets.
In Figure 4, we plot scatter plots between VUA-ratings scores and GPT2-base surprisal from both direct and cloze methods, grouping instances based on the location of the metaphoric word within the sentence.
@@
Figure 4: Scatter plots between metaphor novelty scores of VUA-ratings on the x-axis and surprisal (from GPT2-base) direct (blue) and cloze (red) on the y-axis at different relative positions of metaphoric words in sentences. 0-10% means metaphoric words located at the first 1/10Length1/10*Length of the sentence.